# Suri: Multi-constraint Instruction Following for Long-form Text Generation

### v1

@@ Abstract
Existing research on instruction following largely focuses on tasks with simple instructions and short responses. In this work, we explore multi-constraint instruction following for generating long-form text. We create Suri, a dataset with 20K human-written long-form texts paired with LLM-generated backtranslated instructions that contain multiple complex constraints. Because of prohibitive challenges associated with collecting human preference judgments on long-form texts, preference-tuning algorithms such as DPO are infeasible in our setting; thus, we propose Instructional ORPO (I-ORPO), an alignment method based on the ORPO algorithm. Instead of receiving negative feedback from dispreferred responses, I-ORPO obtains negative feedback from synthetically corrupted instructions generated by an LLM. Using Suri, we perform supervised and I-ORPO fine-tuning on Mistral-7b-Instruct-v0.2. The resulting models, Suri-SFT and Suri-I-ORPO, generate significantly longer texts ($\sim$5K tokens) than base models without significant quality deterioration. Our human evaluation shows that while both SFT and I-ORPO models satisfy most constraints, Suri-I-ORPO generations are generally preferred for their coherent and informative incorporation of the constraints. [^1] Code at https://github.com/chtmp223/suri
Suri: Multi-constraint Instruction Following for Long-form Text Generation
@@ 1 Introduction
We conduct a series of automatic and human evaluations on generations from SFT and I-ORPO-tuned models to validate our method. Compared to the base model, Mistral-7b-Instruct-v0.2 (Jiang et al., 2023), both SFT and I-ORPO significantly increase the generation length from 1K to 5K tokens. Our fine-tuned models also improve the ability to differentiate between correct and corrupted instructions by at least 10% while maintaining low levels of $n$-gram repetitions in the text. We find that GPT-4o (OpenAI, 2024) cannot reliably evaluate long-form responses, making human evaluation crucial for assessing the constraint-following capabilities of our generations. Annotators note that our fine-tuned models effectively follow given constraints, with I-ORPO being preferred for its ability to incorporate constraints coherently, informatively, and enjoyably.
@@ 2 The Suri Dataset
Writing Instructions WritingPrompt (Fan et al., 2018) 300K Creative Writing 28 735
ROCStory (Mostafazadeh et al., 2016) 50K Creative Writing 36 8
@@
Table 1: Comparison of Suri with other single-turn datasets in terms of relevant data features, including writing instructions, instruction-following datasets, and constrained instructions. The data size as well as the average length of prompts and responses (measured in words) is either quoted from the original paper or calculated from publicly available subsets. Suri is the only dataset featuring both constrained instructions and long responses (> 4K words) specifically designed for text generation.
@@ 2.1 Collecting Responses
RedPajama contains over 100 billion documents from 84 CommonCrawl dumps, of which 6,646 texts are sampled. Unlike ChapterBreak and Books3, which consist primarily of books and literary narratives, RedPajama captures the style of everyday writing with informal textual content such as blog posts, obituaries, and more. We apply a set of quality filters (see Appendix A) on the 2023-06 and 2023-14s snapshots to obtain a subset of $\approx 300$K high-quality, non-duplicated documents written in English.
@@ 2.2 Creating Instructions via Backtranslation
Suri includes backtranslated instructions ($x_{w}$) and corrupted instructions ($x_{l}$). In $x_{l}$, constraints from $x_{w}$ are minimally edited to be partially violated while still faithful to the overall main goal of the instruction. These corrupted instructions, along with $x_{w}$ and $y$, serve as inputs for our I-ORPO preference tuning experiments.
@@
Our extracted gold responses do not come with accompanying instructions. Gathering these instructions can be costly and time-consuming, as annotators have to synthesize the instructions from long texts. Therefore, we use instruction backtranslation (Li et al., 2023; Köksal et al., 2023) to generate the missing instructions. Specifically, we prompt GPT-4-turbo [^4] GPT-4-turbo refers to gpt-4-0215-preview. Experiment done using temperature=0.6 and top_p=0.9. with a gold response $y$ to generate a corresponding instruction $x_{w}$ that contains a main goal, which summarizes the content of the text, and a list of $\approx$10 constraints (Table 9). These constraints can focus on stylistic elements (how something is communicated through tone, language, sentence structure), semantic elements (what topics, meanings, and concepts are included), or a combination of both. Constraints can also be broad, applying to large portions of the text, or specific, addressing elements that occur only once. The result is a highly detailed, multi-constraint instruction that covers different parts of the text ($x_{w}$ in Figure 1).
@@
We want to use Suri in our alignment experiments, which traditionally rely on preference judgments (e.g., labeled $y_{w}$ and $y_{l}$ pairs). However, obtaining these judgments for long-form outputs is challenging due to the many competing aspects to consider (e.g., faithfulness to instructions, overall coherence, etc.). Instead of including a corrupted $y_{l}$, we focus on learning from a corrupted instruction $x_{l}$. To create $x_{l}$, we prompt GPT-4-turbo [^5] Experiment done using model=gpt-4-0125-preview, temperature=0.0, top_p=0.0 to ensure deterministic results. to minimally edit each constraint in $x_{w}$ while preserving the original main goal (Table 10). The resulting instructions average 386 tokens, closely matching the average length of gold instructions at approximately 411 tokens.
@@ 2.3 Validating Instructions
To validate whether the backtranslated instructions faithfully represent the original text, we conduct a human evaluation on a sample of 30 $(x_{w},y)$ pairs. Three Upwork [^6] See Appendix F for recruitment and compensation details. annotators are asked to read through the $(x_{w},y)$ pairs, highlight all text spans in the response that support the given constraints, and determine if the response supports the instruction (Figure 6). Our findings indicate that, on average, about 87% of the constraints are fully satisfied, with the remaining constraints being partially satisfied (see Appendix F for agreement statistics). We conclude that the backtranslated instructions are generally faithful to the original text.
@@ 3 Aligning language models with Suri
Odds Ratio Preference Optimization (ORPO) Hong et al. (2024) combines SFT and preference alignment by incorporating a log odds ratio term into the negative log-likelihood loss. We choose ORPO due to its simplicity, competitive performance with other preference tuning algorithms and the ease with which we can modify for our setting. The original algorithm learns from preference judgments, requiring access to chosen and rejected responses in the $(x,y_{w},y_{l})$ format. Since our dataset contains gold and corrupted instructions instead, we modify ORPO so that the algorithm accepts $(x_{w},x_{l},y)$. We refer to this modified method as Instructional Odds Ratio Preference Optimization (I-ORPO), where the modified loss is calculated as:
@@
In the original ORPO formulation, the model is learning if the log probability of $P_{\theta}(y_{w}|x)$, denoted $\text{logps}(y_{w}|x)$, increases and log probability of $P_{\theta}(y_{l}|x)$, denoted $\text{logps}(y_{l}|x_{w})$, decreases after a number of training steps, resulting in the log odds ratio increasing. In I-ORPO, the same $y$ is used for both instruction types. Therefore, the model is learning if the log probabilities $\text{logps}(y|x_{w})$ and $\text{logps}(y|x_{l})$ diverge while $\text{logps}(x_{w})$ and $\text{logps}(x_{l})$ remain stable. We observe this trend in Figure 3. Loss derivation and analysis are in Appendix D.
Figure 3: ORPO training curve. Left figure documents the log probability of the chosen and rejected prompts over 2 epochs. Right figure shows the log probability of the response given the chosen and rejected prompts over 2 epochs. A divergence between $\text{logps}(y|x_{w})$ and $\text{logps}(y|x_{l})$ is observed after 0.5 training epoch.
@@ 4.3 I-ORPO improves ranking accuracy
We use Hugging Face’s Transformers Wolf et al. (2020) to access the probability distribution over vocabulary and measure the impact of instruction specificity on ranking accuracy across five different settings, which are defined by the number of all constraints included (M constraints in total) and the number of those included constraints that are corrupted: (M,M), (M,M/2), (M,1), (M/2,M), (1,M). For example, in the (M, M/2) setting, both instructions include all constraints, but only half of the constraints are violated.
@@
We experiment with using LLMs to evaluate whether texts generated by our fine-tuned models follow the given constraints. Specifically, we provide GPT-4o (OpenAI, 2024) with a constraint and a generated text from our models and prompt it to determine whether the text fully satisfies, partially satisfies, or does not satisfy the constraint (Table 16). We then compare these results with judgments from three Upwork annotators on 30 texts generated by Suri-SFT on the test set (obtained using the same procedure as in Section 2.3). GPT-4o agrees with human annotators only 39% of the time, with a significant 16% disagreement between satisfaction and no satisfaction (Table 4). We conclude that GPT-4o does not align well with long-form human annotation, consistent with the findings of Xu et al. (2024) and Kim et al. (2024).
GPT vs Human Percentage
Agreement 39%
Partial satisfaction - No satisfaction 23%
Satisfaction - Partial satisfaction 22%
Satisfaction - No satisfaction 16%
Table 4: Types of disagreement between GPT-4o and human judges on 30 samples generated from Suri-SFT.
@@ 5 Human Evaluation
While our automatic assessments provide insights into the lexical information of the text, they do not capture its semantic content. Therefore, we conduct a human evaluation to determine if and how the constraints are satisfied by the outputs of Suri-SFT and Suri-I-ORPO. Human evaluation on 30 test set generations reveals that while both fine-tuned models satisfy constraints, Suri-I-ORPO is preferred by humans for its ability to incorporate the constraints into the final outputs seamlessly.
@@
Since GPT-4o judgments do not align with human annotations, we rely on human evaluation to determine how often Suri-I-ORPO and Suri-SFT follow the given constraints. This evaluation follows a similar setup as Section 2.3, where annotators assess whether each constraint is satisfied, partially satisfied, or not satisfied by the generations. Two Upwork annotators complete 30 tasks, each containing a generation with around 10 constraints, totaling 321 constraints. The generations are lengthy, averaging 4,000 words, and complex, with constraints spread throughout the text. Annotators spend approximately 20-25 minutes on each annotation and are paid \$200 in total for the task.
@@ 7 Conclusion
In this work, we investigate the challenge of complex instruction following for generating long-form text. We introduce Suri, a dataset of long human-written responses accompanied by backtranslated and corrupted instructions. We demonstrate the effectiveness of Suri in improving the constraint-following capabilities of LLMs for long-form generation through supervised fine-tuning and I-ORPO. Our models, as shown by both human and automated evaluations, generate high-quality, long-form responses while maintaining effectiveness at following constraints.
@@
Extending to other LLMs
While we demonstrate the effectiveness of Suri and I-ORPO on Mistral-7b-Instruct-v0.2, we have yet to experiment with fine-tuning other models on our dataset, which presents an interesting direction for future work.
@@
Even though I-ORPO works well on our dataset, we would like to explore how surface features, such as instruction length and the degree of information overlap between the instruction and response, affect its performance. We leave this investigation to future studies.
Performance on short-context tasks
Additionally, we note that our dataset primarily focuses on generating extremely long texts. As a result, the fine-tuned models may exhibit diminished performance on tasks requiring shorter generations.
@@
The risks posed by our study are no greater than those inherent in the large language models that support it (Weidinger et al., 2021). Our human evaluation receives approval from an institutional review board. All annotators (US-based, fluent in English) gave their informed consent and participated with an hourly compensation of \$16, which meets the minimum wage in our state. Scientific artifacts are implemented according to their intended usage.
@@ Acknowledgements
We extend special gratitude to the Upwork annotators for their hard work, as well as to the members of Unsloth, r/LocalLLaMA, and Together.ai community for helpful fine-tuning advice. We also thank Scott Niekum, Dzung Pham, and members of the UMass NLP lab for their insights on the project. This project was partially supported by awards IIS-2202506 and IIS-2312949 from the National Science Foundation (NSF) and an award from Open Philanthropy.

### v2

@@ Abstract
Existing research on instruction following largely focuses on tasks with simple instructions and short responses. In this work, we explore multi-constraint instruction following for generating long-form text. We create Suri, a dataset with 20K human-written long-form texts paired with LLM-generated backtranslated instructions that contain multiple complex constraints. Because of prohibitive challenges associated with collecting human preference judgments on long-form texts, preference-tuning algorithms such as DPO are infeasible in our setting; thus, we propose Instructional ORPO (I-ORPO), an alignment method based on the ORPO algorithm. Instead of receiving negative feedback from dispreferred responses, I-ORPO obtains negative feedback from synthetically corrupted instructions generated by an LLM. Using Suri, we perform supervised and I-ORPO fine-tuning on Mistral-7b-Instruct-v0.2. The resulting models, Suri-SFT and Suri-I-ORPO, generate significantly longer texts ($\sim$5K tokens) than base models without significant quality deterioration. Our human evaluation shows that while both SFT and I-ORPO models satisfy most constraints, Suri-I-ORPO generations are generally preferred for their coherent and informative incorporation of the constraints. [^1] Code & Data are available at https://github.com/chtmp223/suri
Suri: Multi-constraint Instruction Following for
Long-form Text Generation
@@ 1 Introduction
We conduct a series of automatic and human evaluations on generations from SFT and I-ORPO-tuned models to validate our method. Compared to the base model, Mistral-7b-Instruct-v0.2 (Jiang et al., 2023), both SFT and I-ORPO significantly increase the generation length from 1K to 5K tokens. Our fine-tuned models also improve the ability to differentiate between correct and corrupted instructions by at least 10% while maintaining low levels of $n$-gram repetitions in the text. We find that LLM judges, such as GPT-4o (OpenAI, 2024), Gemini-1.5-Pro (Team et al., 2024), and Claude-3.5-Sonnet (Anthropic, 2024), cannot reliably evaluate long-form responses, which makes human evaluation crucial for assessing the constraint-following capabilities of our generations. Annotators note that our fine-tuned models effectively follow given constraints, with I-ORPO being preferred to our SFT model for its ability to incorporate constraints coherently, informatively, and enjoyably.
@@ 2 The Suri Dataset
Writing Instructions ROCStory (Mostafazadeh et al., 2016) 50K Creative Writing 36 8
WritingPrompt (Fan et al., 2018) 300K Creative Writing 28 735
@@
Table 1: Comparison of Suri with other single-turn datasets in terms of relevant data features, including writing instructions, instruction-following datasets, and constrained instructions. The data size as well as the average length of prompts and responses is either quoted from the original paper or calculated from publicly available subsets. Suri is the only dataset featuring both constrained instructions and long responses (> 4K tokens) specifically designed for text generation.
@@ 2.1 Collecting Responses
RedPajama contains over 100 billion documents from 84 CommonCrawl dumps. Unlike ChapterBreak and Books3, which consist primarily of books and literary narratives, RedPajama captures the style of everyday writing with informal textual content such as blog posts, obituaries, and more. We apply a set of quality filters (see Appendix A) on the 2023-06 and 2023-14 snapshots to obtain a subset of $\approx 300$K high-quality, non-duplicated documents written in English, from which 6,646 texts are sampled.
@@ 2.2 Creating Instructions via Backtranslation
Suri includes backtranslated instructions ($x_{w}$) and corrupted instructions ($x_{l}$). To create $x_{l}$, constraints from $x_{w}$ are minimally edited to be partially violated while still faithful to the overall main goal of the instruction. These corrupted instructions $x_{l}$, along with $x_{w}$ and $y$, serve as inputs for our I-ORPO preference tuning experiments.
@@
Our extracted gold responses do not come with accompanying instructions. Gathering these instructions can be costly and time-consuming, as annotators have to synthesize the instructions from long texts. Therefore, we use instruction backtranslation (Li et al., 2023; Köksal et al., 2023) to generate the missing instructions. Specifically, we prompt GPT-4-turbo [^4] GPT-4-turbo refers to gpt-4-0215-preview. Experiment done using temperature=0.6 and top_p=0.9. with a gold response $y$ to generate a corresponding instruction $x_{w}$ that contains a main goal, which summarizes the content of the text, and a list of $\approx$10 constraints (Table 9). These constraints can focus on stylistic elements (how something is communicated through tone, language, sentence structure), semantic elements (what topics, meanings, and concepts are included), or a combination of both. Constraints can also be broad, applying to large portions of the text, or specific, addressing elements that occur only once. The result is a set of highly detailed, multi-constraint instructions that cover different parts of the text ($x_{w}$ in Figure 1).
@@
We want to use Suri in our alignment experiments, which traditionally rely on preference judgments (e.g., labeled $y_{w}$ and $y_{l}$ pairs). However, obtaining these judgments for long-form outputs is challenging due to the many competing aspects to consider (e.g., faithfulness to instructions, overall coherence, etc.). Therefore, we focus on learning from a corrupted instruction $x_{l}$ instead of from $y_{l}$. To create $x_{l}$, we prompt GPT-4-turbo [^5] Experiment done using model=gpt-4-0125-preview, temperature=0.0, top_p=0.0 to ensure deterministic results. to minimally edit each constraint in $x_{w}$ while preserving the original main goal (Table 10). The resulting instructions average 386 tokens, closely matching the average length of gold instructions at approximately 411 tokens. [^6] The first author manually reviews a random subset of 50 corrupted claims by comparing them to their corresponding original versions, and confirms that the corruptions are minimal and effective in invalidating the original constraints.
@@ 2.3 Validating Instructions
To validate whether the backtranslated instructions faithfully represent the original text, we conduct a human evaluation on a sample of 30 $(x_{w},y)$ pairs. Three Upwork [^7] See Appendix F for recruitment and compensation details. annotators are asked to read through the $(x_{w},y)$ pairs, highlight all text spans in the response that support the given constraints, and determine if the response supports the instruction (Figure 6).
Our findings indicate that, on average, about 87% of the constraints are fully satisfied, with the remaining constraints being partially satisfied (see Appendix F for agreement statistics). We conclude that the backtranslated instructions are generally faithful to the original text.
@@ 3 Aligning language models with Suri
Odds Ratio Preference Optimization (ORPO) Hong et al. (2024) combines SFT and preference alignment by incorporating a log odds ratio term into the negative log-likelihood loss. We choose ORPO due to its simplicity, competitive performance with other preference tuning algorithms and the ease with which we can modify for our setting.
The original algorithm learns from preference judgments, requiring access to chosen and rejected responses in the $(x,y_{w},y_{l})$ format. Since our dataset contains gold and corrupted instructions instead, we modify ORPO so that the algorithm accepts $(x_{w},x_{l},y)$. We refer to this modified method as Instructional Odds Ratio Preference Optimization (I-ORPO), where the modified loss is calculated as:
@@
In the original ORPO formulation, the model is learning if the log probability of $P_{\theta}(y_{w}|x)$, denoted $\text{logps}(y_{w}|x)$, increases and log probability of $P_{\theta}(y_{l}|x)$, denoted $\text{logps}(y_{l}|x_{w})$, decreases after a number of training steps, resulting in the log odds ratio increasing. In I-ORPO, the same $y$ is used for both instruction types. Therefore, the model is learning if the log probabilities $\text{logps}(y|x_{w})$ and $\text{logps}(y|x_{l})$ diverge while $\text{logps}(x_{w})$ and $\text{logps}(x_{l})$
remain stable. We observe this trend in Figure 3. Loss derivation and analysis are in Appendix D.
Figure 3: ORPO training curve. Left figure documents the log probability of the chosen and rejected prompts over 3 epochs. Right figure shows the log probability of the response given the chosen and rejected prompts over 3 epochs. A divergence between $\text{logps}(y|x_{w})$ and $\text{logps}(y|x_{l})$ is observed after 0.5 training epoch.
@@
For both Suri-SFT and Suri-I-ORPO, we set the number of epochs to 2, which is determined by a manual inspection of model generations at each epoch checkpoint. We observe that as the number of epochs increases from 1 to 5, both the response length and the number of 5-gram repetitions increase, indicating a trade-off between response length and repetition. After reviewing 30 generations at each epoch, we select the configuration that produces the most coherent responses and minimal repetition. Based on these criteria, the optimal number of epochs is 2, which balances the trade-off between response length and response quality.
@@ 4.3 I-ORPO improves ranking accuracy
We use Hugging Face’s Transformers Wolf et al. (2020) to access the probability distribution over vocabulary and measure the impact of instruction specificity on ranking accuracy across five different settings, which are defined by the number of all constraints included (M constraints in total) and the number of those included constraints that are corrupted: (M,M), (M,M/2), (M,1), (M/2,M/2), (1,1). For example, in the (M, M/2) setting, both instructions include all constraints, but only half of the constraints are violated.
@@
We experiment with using LLMs to evaluate whether texts generated by our fine-tuned models follow the given constraints. Specifically, we provide GPT-4o (OpenAI, 2024), Gemini-1.5-Pro (Team et al., 2024), and Claude-3.5-Sonnet (Anthropic, 2024) with a constraint and a generated text from our models and prompt it to determine whether the text fully satisfies, partially satisfies, or does not satisfy the constraint (Table 16). [^12] Experiment done using OpenAI API for GPT-4o and Vertex API for Claude-3.5-Sonnet and Gemini-1.5-Pro. Temperature is set to 0.0 and the maximum number of generated tokens is set to 4096 for all models. We then compare these results with judgments from three Upwork annotators on 30 texts generated by Suri-SFT on the test set (obtained using the same procedure as in Section 2.3). GPT-4o agrees with human annotators only 39% of the time, with a significant 16% disagreement between satisfaction and no satisfaction (Table 4). Claude-3.5-Sonnet and Gemini-1.5-Pro lagging significantly behind, with Claude agreeing with humans only 24% of the time, and Gemini 13% of the time. Notably, Gemini refuses to annotate 39% of the time, even when all safety filters have been disabled. We conclude that LLM judges do not align well with long-form human annotation, consistent with the findings of Xu et al. (2024) and Kim et al. (2024).
Category GPT-4o Claude-3.5-Sonnet Gemini-1.5-Pro
Agreement with human’s majority vote 39% 24% 13%
Partial satisfaction - No satisfaction 23% 6% 2%
Satisfaction - Partial satisfaction 22% 63% 1%
Satisfaction - No satisfaction 16% 7% 45%
Table 4: Types of agreement and disagreement between GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and human judges on 30 generations from Suri-SFT.
@@ 5 Human Evaluation
While our automatic assessments provide insights into the lexical information of the text, they do not capture its semantic content. Therefore, we conduct a human evaluation to determine if and how the constraints are satisfied by the outputs of Suri-SFT and Suri-I-ORPO. Human evaluation on 30 test set generations reveals that while both fine-tuned models satisfy constraints, Suri-I-ORPO is preferred by humans for its ability to seamlessly incorporate the constraints into the final outputs.
@@
Since GPT-4o judgments do not align with human annotations, we rely on human evaluation to determine how often Suri-I-ORPO and Suri-SFT follow the given constraints. This evaluation follows a similar setup as Section 2.3, where annotators assess whether each constraint is satisfied, partially satisfied, or not satisfied by the generations. Two Upwork annotators complete 30 tasks, each containing a generation with around ten constraints, totaling 321 constraints. The generations are lengthy, averaging 4,000 words, and complex, with constraints spread throughout the text. Annotators spend approximately 20-25 minutes on each annotation and are paid \$200 in total for the task.
@@ 7 Conclusion
In this work, we investigate the challenge of complex instruction following for generating long-form text. We introduce Suri, a dataset of long human-written responses accompanied by backtranslated and corrupted instructions. We demonstrate the effectiveness of Suri in improving the constraint-following capabilities of LLMs for long-form generation through supervised fine-tuning and I-ORPO. Human and automated evaluations show that our models generate high-quality, long-form responses while effectively satisfying constraints.
@@
Fine-tuning additional LLMs on Suri
While we demonstrate the effectiveness of Suri and I-ORPO on Mistral-7b-Instruct-v0.2, we have yet to experiment with fine-tuning other models on our dataset using I-ORPO.
@@
Even though I-ORPO works well on our dataset, we would like to explore how surface features, such as instruction length and the degree of information overlap between the instruction and response, affect its performance.
Impact of truncating gold responses
In our experiments, we truncate gold responses to lengths between 2,048 and 5,024 words to make fine-tuning more cost-effective and computationally efficient. However, our released code includes an option that allows users to recover the full response text, and thus bypass the truncated version if needed.
Ranking accuracy on out-of-domain datasets
We report the ranking accuracy on the Suritest set, where Suri-SFT and Suri-I-ORPO may have an advantage over the baseline models due to their fine-tuning on Suri.
@@
Our human evaluation receives approval from an institutional review board. All annotators (US-based, fluent in English) gave their informed consent and participated with an hourly compensation of \$16, which meets the minimum wage in our state. Scientific artifacts are implemented according to their intended usage.
@@ Acknowledgements
We extend special gratitude to the Upwork annotators for their hard work and to the members of Unsloth, r/LocalLLaMA, and Together.ai community for helpful fine-tuning advice. We also thank Scott Niekum, Dzung Pham, and members of the UMass NLP lab for their insights on the project. This project was partially supported by awards IIS-2202506 and IIS-2312949 from the National Science Foundation (NSF) and an award from Open Philanthropy.